Find the neutral temperature and inversion temperature of copper-iron thermocouple if the reference junction is kept at 0°C. Use the data in table (33.1).

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. θ = 0°C a base Cu,Fe = a base Cu,Pb – a base Fe,Pb = 2.76 – 16.6 = – 13.8 μv/°C BCu,Fe = bCu,Pb – bFe,Pb = 0.012 + 0.030 = 0.042 μv/°C^2 Neutral temp. on – a/b = 13.8/0.042 °C = 328.57°C